# Duncan McCoy and Connor Sell

### v2

@@ 5. Rational quadratic forms
We begin by recalling some facts concerning the theory of rational quadratic forms and Hilbert symbols. For a detailed account we refer the reader to [Ser73]. For each prime $p$ we have the Hilbert symbol (_,_$(\_,\_)_{p}:\mathbb{Q}_{p}^{\times}\times\mathbb{Q}_{p}^{\times}\rightarrow\{\pm 1\}$, which is defined to be
$(a,b)_{p}=\begin{cases}1&\text{if $z^{2}=ax^{2}+by^{2}$ has a nonzero solutionin $\mathbb{Q}_{p}$ }\\ -1&\text{otherwise.}\end{cases}$
@@
By [Ser73, Ch. III, Thm. 4], there exists a $c\in\mathbb{Q}^{\times}$ such that $(d,c)_{p}=h_{p}$ for all $p$. Since, there are an even number of $p$ such that $h_{p}=-1$, we have that $(c,d)_{\infty}=1$. Thus, if $d<0$, then $c$ is automatically positive. If $d>0$ and $c<0$, then the observation that $(c,d)_{p}=(c,d)_{p}(-d,d)_{p}=(-cd,d)_{p}$ for all primes $p$ shows that $-cd>0$ satisfies the necessary conclusions. ∎
@@ 5.2. Projective equivalence
We will make use of the following weaker notion of equivalence for quadratic forms.
Definition 5.3 (Projective equivalence).
@@
It is clear that the signature of a form $q$ is invariant under scaling by a positive number and that the discriminant of a rank $n$ form satisfies $d(mq)=m^{n}d(q)$. Since the discriminant lies in ×/(×$\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$, it is invariant under projective equivalence if and only if the rank $n$ of $q$ is even; if $n$ is odd, any projective equivalence class contains quadratic forms of all possible discriminants. The Hasse-Witt invariants of a scaled form depend on the value of $n$ (mod 4).
@@
$\epsilon_{p}(mq)=\begin{cases}\epsilon_{p}(q)&\text{if $n\equiv 1\bmod 4$}\\ \epsilon_{p}(q)(m,-d(q))_{p}&\text{if $n\equiv 2\bmod 4$}\\ \epsilon_{p}(q)(m,-1)_{p}&\text{if $n\equiv 3\bmod 4$}\\ \epsilon_{p}(q)(m,d(q))_{p}&\text{if $n\equiv 0\bmod 4$}\\ \end{cases}$
@@ 5.3. Combining forms
Theorem 5.2 implies that there exist rationals $a^{\prime},b^{\prime},c^{\prime}\in\mathbb{Q}^{\times}$ such that the form $\langle a^{\prime},b^{\prime},c^{\prime}\rangle$ has Hasse-Witt invariants $\epsilon_{p}=h^{\prime}_{p}$ for every $p$, discriminant $dxyz$ and the same signature as $\langle x,y,z\rangle$. Since $\langle a^{\prime},b^{\prime},c^{\prime}\rangle$ and $\langle x,y,z\rangle$ have the signature, we can assume that the quantities $a=a^{\prime}x$, $b=b^{\prime}y$, and $c=c^{\prime}z$ are all positive. By construction, we have that
@@ 6.2. Real representations
where $q_{i}\in\mathcal{Q}(\rho_{i})$ for each $i$. Clearly $q$ is non-degenerate if and only if each of the $q_{i}$ is non-degenerate. ∎
@@ 6.3. Non-real representations
Let $q$ be a non-degenerate quadratic form of even rank $2n$ over a field of characteristic 00. If $q$ has an $n$-dimensional isotropic subspace, then $q$ is equivalent to the form
@@
where $Q^{\prime}$ represents a form in $\mathcal{Q}(\rho^{m-2})$. This allows us to proof the lemma by induction on $m$. ∎
@@
It is a theorem of Brauer that the cyclotomic field $\mathbb{Q}[\zeta_{e}]$ is a splitting field for $G$ [Ser77, Theorem 24]. This means that any representation of $G$ that is irreducible over $\mathbb{Q}[\zeta_{e}]$ remains irreducible over any field extension of $\mathbb{Q}[\zeta_{e}]$. In particular, if we decompose $\rho$ into irreducible representations over $\mathbb{Q}[\zeta_{e}]$, then these representations remain irreducible over $\mathbb{C}$ and so are all non-real by hypothesis. By Lemma 6.3, Lemma 6.8 and Lemma 6.9, this means that the form $q$ is equivalent to $\langle 1,-1\rangle^{\oplus\frac{n}{2}}$ over $\mathbb{Q}[\zeta_{e}]$. For $p>2$, the $p$-adic field $\mathbb{Q}_{p}$ contains $(p-1)$-roots of unity [Ser73, Ch. II, Prop. 7]. Therefore, for $p\equiv 1\bmod e$ the $p$-adic field $\mathbb{Q}_{p}$ contains a copy of $\mathbb{Q}[\zeta_{e}]$. Thus $q$ is equivalent $\langle 1,-1\rangle^{\oplus\frac{n}{2}}$ over such a $\mathbb{Q}_{p}$, allowing us to compute the Hasse-Witt invariants and discriminant of $q$ in this field. ∎
@@ 7.2. Manifolds with odd holonomy
On the other hand it is possible to use Theorem 1.1 to obstruct for certain flat manifolds from occurring as a cusp cross-section in some commensurability classes. Notably we are able to obtain obstructions in all dimensions.
@@
Now fix a prime $p\equiv 1\bmod 4e$. Since $p\equiv 1\bmod 4$, we have that $-1$ is a square in $\mathbb{Q}_{p}$. Since $p\equiv 1\bmod e$, Proposition 6.10 implies that $d(g)=1\in\mathbb{Q}_{p}^{\times}/(\mathbb{Q}_{p}^{\times})^{2}$ and $\epsilon_{p}(g)=1$. In particular this implies that $q$ is equivalent to $\langle a_{1},\dots,a_{m},1,\dots,1\rangle$ over $\mathbb{Q}_{p}$. This allows us to easily calculate $d(q)\in\mathbb{Q}_{p}^{\times}/(\mathbb{Q}_{p}^{\times})^{2}$ and $\epsilon_{p}(q)$.
@@
If $b_{1}(B)\geq 3$, then Corollary 7.2 shows that $B$ appears as a cusp cross-section in every commensurability class of cusped arithmetic hyperbolic $(n+1)$-manifolds. Conversely, let $q$ be a quadratic form of signature $(n+1,1)$ with $d(q)=-1$ and $\epsilon_{p}(q)=-1$ for some prime $p\equiv 1\bmod 4|H|$, where $H$ is the holonomy group of $B$. If $b_{1}(B)\leq 2$, then Theorem 7.4 shows that $B$ does not arise as a cusp cross-section in the commensurability class defined by $q$. Since $d(q)=-1$, Proposition 5.5 implies that $\epsilon_{p}(q)$ at such a prime is an invariant of the projective equivalence class of $q$. Since there are infinitely many primes satisfying $p\equiv 1\bmod 4|H|$, we may construct infinitely many distinct commensurability classes which do not contain $B$ as a cusp cross-section. ∎
@@ 7.3. The flat 3- and 4-manifolds
We first recover the classification of cusp cross-sections for hyperbolic arithmetic 4-manifolds as derived by the second author [Sel23]. The compact orientable flat 3-manifolds are listed in Table 1. Note that for five dimensional quadratic forms the projective invariants are simply the Hasse-Witt invariants $\epsilon_{p}(q)$ (see Proposition 5.5).
@@
Since $-3$ is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\mod 3$, this implies that $\epsilon_{p}(q)=(3,-1)_{p}$ whenever $p\equiv 1\bmod 3$. Since $(3,-1)_{p}=1$ unless $p=2$ or 3, this implies that $\epsilon_{p}(q)=1$ whenever $p\equiv 1\bmod 3$. Conversely, every form $q$ of signature $(4,1)$ satisfying $\epsilon_{p}(q)=1$ for all $p\equiv 1\mod 3$ is equivalent to $\langle 3a,a,b,1,-1\rangle$ for some $a,b\in\mathbb{Q}_{>0}$. Proposition 5.1 gives $c>0$ such that $(-3,c)_{p}=\epsilon_{p}(q)(3,-1)_{p}$ all primes. Setting $b=-3d(q)$ and $a=-3d(q)c$ yields a form with equivalent to $q$.
@@
Since $-1$ is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\mod 4$, this implies that $\epsilon_{p}(q)=1$ whenever $p\equiv 1\bmod 4$.
@@
3-torus/$O^{3}_{1}$ $C_{1}$ $3$ None
$\frac{1}{2}$-twist/$O^{3}_{2}$ $C_{2}$ $1$ None
$\frac{1}{3}$-twist/$O^{3}_{3}$ $C_{3}$ $1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod{3}$
$\frac{1}{4}$-twist/$O^{3}_{4}$ $C_{4}$ $1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod{4}$
$\frac{1}{6}$-twist/$O^{3}_{5}$ $C_{6}$ $1$ $\epsilon_{p}(q)=1$ for $p\equiv 1\bmod{3}$
Hantzsche-Wendt/$O^{3}_{6}$ $C_{2}^{2}$ 00 None
@@
For quadratic forms of rank six, Proposition 5.5 shows that invariants of the projective equivalence class are the discriminant $d(q)$ and $\epsilon_{p}(q)$ for all primes $p$ such that $-d(q)$ is a square in $\mathbb{Q}_{p}$.
@@
    $H$ is $C_{4}$ or $D_{8}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\bmod 4$ and $-d(q)$ is a square in $\mathbb{Q}_{p}$,
@@
    $H$ is $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\mod 3$ and $-d(q)$ is a square in $\mathbb{Q}_{p}$, or
@@
    $H$ is $A_{4}$, and $\epsilon_{p}(q)=1$ for all primes $p$ such that $-d(q)$ is a square in $\mathbb{Q}_{p}$.
@@
  • If $H$ is isomorphic to $C_{4}$ or $D_{8}$, the the representation $\rho_{1}$ must contain an element of order 4. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent to a quadratic form $\langle a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies $d(q)=-bc$ and one can calculate its Hasse-Witt invariants to be
@@
    Thus, we see that $\epsilon_{p}(q)=1$ whenever both $-d(q)$ and $-1$ are squares in $\mathbb{Q}_{p}$. However $-1$ is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 4$.
    Conversely, any form $q$ of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ for all $p$ such that $-d(q)$ and $-1$ are both squares in $\mathbb{Q}_{p}$ is equivalent to one of the form $\langle a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Proposition 5.1 allows us to pick $\alpha\in\mathbb{Q}_{>0}$ such that $(\alpha,-1)_{p}=\epsilon_{p}(q)$ for all primes where $-d(q)$ is a square in $\mathbb{Q}_{p}$. Thus we see that $\epsilon_{p}(q)(\alpha,-1)_{p}=1$ for all primes where $-d(q)$ is a square in $\mathbb{Q}_{p}$. This allows us to apply Proposition 5.1 again to obtain $\beta\in\mathbb{Q}_{>0}$ such that $(\beta,-d(q))_{p}=\epsilon_{p}(q)(\alpha,-1)_{p}$ for all primes. Consequently we have $(\beta d(q),-d(q))_{p}=\epsilon_{p}(q)(\alpha,-1)_{p}$ for all primes. If we set $b=-\beta d(q)>0$, $a=\alpha b$ and $c=\beta$, then this yields a form equivalent to $q$.
  • If $H$ is isomorphic to $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$, then $\rho_{1}$ must contain an element of order 3. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent to $\langle 3a,a,b,c,1,-1\rangle$, for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies $d(q)=-3bc$ and
@@
    Thus, we see that $\epsilon_{p}(q)=(3,-1)_{p}$ whenever both $-d(q)$ and $-3$ are squares in $\mathbb{Q}_{p}$. However $-3$ is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 3$. Since $(3,3)_{p}=-1$ if and only if $p=2,3$, this reduces to the condition that $\epsilon_{p}(q)=1$.
    As in the previous case a double application of Proposition 5.1 can be used to show that any form $q$ of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ for all $p$ such that $-d(q)$ and $-3$ are both squares in $\mathbb{Q}_{p}$ is equivalent to $\langle 3a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$.
Thus it remains to treat the case that $H$ is the alternating group $A_{4}$. Since we know the holonomy representation is faithful, we see that $\rho$ must contain a copy of the unique irreducible 3-dimensional representation of $A_{4}$. Thus we the holonomy representation decomposes as $\rho\cong\rho_{1}\oplus\rho_{2}$, where $\rho_{1}$ is 3-dimensional and $\rho_{2}$ is 1-dimensional. By direct calculation one can show that every form in $\mathcal{Q}(\rho_{1})$ is equivalent to $\langle a,a,a\rangle$ for some $a\in\mathbb{Q}$. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent $\langle a,a,a,b,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies $d(q)=-ab$ and
@@
Thus we have that $\epsilon_{p}(q)=1$ whenever $-d(q)$ is a square in $\mathbb{Q}_{p}$. Once again, Proposition 5.1 can be used to obtain the converse. If $q$ is a quadratic form of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ whenever $-d(q)$ is a square in $\mathbb{Q}_{p}$, then there is $b\in\mathbb{Q}_{>0}$ such that $\epsilon_{p}(q)=(b,-d(q))_{p}=(bd(q),-d(q))_{p}$ for all $p$. If we set $a=-bd(q)$, then $q$ is equivalent to $\langle a,a,a,b,1,-1\rangle$. ∎

### v3

@@ 5. Rational quadratic forms
We begin by recalling some facts concerning the theory of rational quadratic forms and Hilbert symbols. For a detailed account we refer the reader to [Ser73]. For each prime $p$ we have the Hilbert symbol $(\_,\_)_{p}:\mathbb{Q}_{p}^{\times}\times\mathbb{Q}_{p}^{\times}\rightarrow\{\pm 1\}$, which is defined to be
[…]
@@
By [Ser73, Ch. III, Thm. 4], there exists a $c\in\mathbb{Q}^{\times}$ such that $(d,c)_{p}=h_{p}$ for all $p$. Since there are an even number of $p$ such that $h_{p}=-1$, we have that $(c,d)_{\infty}=1$. Thus, if $d<0$, then $c$ is automatically positive. If $d>0$ and $c<0$, then the observation that $(c,d)_{p}=(c,d)_{p}(-d,d)_{p}=(-cd,d)_{p}$ for all primes $p$ shows that cd>0-cd>0 satisfies the necessary conclusions. ∎
@@ 5.2. Projective equivalence
It is clear that the signature of a form $q$ is invariant under scaling by a positive number and that the discriminant of a rank $n$ form satisfies $d(mq)=m^{n}d(q)$. Since the discriminant lies in $\mathbb{Q}^{\times}/(\mathbb{Q}^{\times})^{2}$, it is invariant under projective equivalence if and only if the rank $n$ of $q$ is even; if $n$ is odd, any projective equivalence class contains quadratic forms of all possible discriminants. The Hasse-Witt invariants of a scaled form depend on the value of $n$ (mod 4).
@@
[…]
@@ 5.3. Combining forms
Theorem 5.2 implies that there exist rationals $a^{\prime},b^{\prime},c^{\prime}\in\mathbb{Q}^{\times}$ such that the form $\langle a^{\prime},b^{\prime},c^{\prime}\rangle$ has Hasse-Witt invariants $\epsilon_{p}=h^{\prime}_{p}$ for every $p$, discriminant dxyzdxyz and the same signature as $\langle x,y,z\rangle$. Since $\langle a^{\prime},b^{\prime},c^{\prime}\rangle$ and $\langle x,y,z\rangle$ have the same signature, we can assume that the quantities $a=a^{\prime}x$, $b=b^{\prime}y$, and $c=c^{\prime}z$ are all positive. By construction, we have that
@@ 6.2. Real representations
where $q_{i}\in\mathcal{Q}(\rho_{i})$ for each ii. Clearly $q$ is non-degenerate if and only if each of the $q_{i}$ is non-degenerate. ∎
@@ 6.3. Non-real representations
Let $q$ be a non-degenerate quadratic form of rank $2n$ over a field of characteristic 0. If $q$ has an $n$-dimensional isotropic subspace, then $q$ is equivalent to the form
@@
where $Q^{\prime}$ represents a form in $\mathcal{Q}(\rho^{m-2})$. This allows us to prove the lemma by induction on $m$. ∎
@@
It is a theorem of Brauer that the cyclotomic field $\mathbb{Q}[\zeta_{e}]$ is a splitting field for $G$ [Ser77, Theorem 24]. This means that any representation of $G$ that is irreducible over $\mathbb{Q}[\zeta_{e}]$ remains irreducible over any field extension of $\mathbb{Q}[\zeta_{e}]$. In particular, if we decompose $\rho$ into irreducible representations over $\mathbb{Q}[\zeta_{e}]$, then these representations remain irreducible over $\mathbb{C}$ and so are all non-real by hypothesis. By Lemma 6.3, Lemma 6.8 and Lemma 6.9, this means that the form $q$ is equivalent to $\langle 1,-1\rangle^{\oplus\frac{n}{2}}$ over $\mathbb{Q}[\zeta_{e}]$. For $p>2$, the $p$-adic field $\mathbb{Q}_{p}$ contains (p1)(p-1)-roots of unity [Ser73, Ch. II, Prop. 7]. Therefore, for $p\equiv 1\bmod e$ the $p$-adic field $\mathbb{Q}_{p}$ contains a copy of $\mathbb{Q}[\zeta_{e}]$. Thus $q$ is equivalent $\langle 1,-1\rangle^{\oplus\frac{n}{2}}$ over such a $\mathbb{Q}_{p}$, allowing us to compute the Hasse-Witt invariants and discriminant of $q$ in this field. ∎
@@ 7.2. Manifolds with odd holonomy
On the other hand it is possible to use Theorem 1.1 to obstruct certain flat manifolds from occurring as a cusp cross-section in some commensurability classes. Notably we are able to obtain obstructions in all dimensions.
@@
Now fix a prime $p\equiv 1\bmod 4e$. Since $p\equiv 1\bmod 4$, we have that 1-1 is a square in $\mathbb{Q}_{p}$. Since $p\equiv 1\bmod e$, Proposition 6.10 implies that $d(g)=1\in\mathbb{Q}_{p}^{\times}/(\mathbb{Q}_{p}^{\times})^{2}$ and $\epsilon_{p}(g)=1$. In particular this implies that $q$ is equivalent to $\langle a_{1},\dots,a_{m},1,\dots,1\rangle$ over $\mathbb{Q}_{p}$. This allows us to easily calculate $d(q)\in\mathbb{Q}_{p}^{\times}/(\mathbb{Q}_{p}^{\times})^{2}$ and $\epsilon_{p}(q)$.
@@
If $b_{1}(B)\geq 3$, then Corollary 7.2 shows that $B$ appears as a cusp cross-section in every commensurability class of cusped arithmetic hyperbolic $(n+1)$-manifolds. Conversely, let $q$ be a quadratic form of signature $(n+1,1)$ with d(q)=1d(q)=-1 and $\epsilon_{p}(q)=-1$ for some prime $p\equiv 1\bmod 4|H|$, where $H$ is the holonomy group of $B$. If $b_{1}(B)\leq 2$, then Theorem 7.4 shows that $B$ does not arise as a cusp cross-section in the commensurability class defined by $q$. Since d(q)=1d(q)=-1, Proposition 5.4 implies that $\epsilon_{p}(q)$ at such a prime is an invariant of the projective equivalence class of $q$. Since there are infinitely many primes satisfying $p\equiv 1\bmod 4|H|$, we may construct infinitely many distinct commensurability classes which do not contain $B$ as a cusp cross-section. ∎
@@ 7.3. The flat 3- and 4-manifolds
We first recover the classification of cusp cross-sections for hyperbolic arithmetic 4-manifolds as derived by the second author [Sel23]. The compact orientable flat 3-manifolds are listed in Table 1. Note that for five-dimensional quadratic forms the projective invariants are simply the Hasse-Witt invariants $\epsilon_{p}(q)$ (see Proposition 5.4).
@@
Since 3-3 is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\mod 3$, this implies that $\epsilon_{p}(q)=(3,-1)_{p}$ whenever $p\equiv 1\bmod 3$. Since $(3,-1)_{p}=1$ unless $p=2$ or 3, this implies that $\epsilon_{p}(q)=1$ whenever $p\equiv 1\bmod 3$. Conversely, every form $q$ of signature $(4,1)$ satisfying $\epsilon_{p}(q)=1$ for all $p\equiv 1\mod 3$ is equivalent to $\langle 3a,a,b,1,-1\rangle$ for some $a,b\in\mathbb{Q}_{>0}$. Proposition 5.1 gives $c>0$ such that $(-3,c)_{p}=\epsilon_{p}(q)(3,-1)_{p}$ for all primes $p$. Setting b=3d(q)b=-3d(q) and a=3d(q)ca=-3d(q)c yields a form with equivalent to $q$.
@@
Since 1-1 is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\mod 4$, this implies that $\epsilon_{p}(q)=1$ whenever $p\equiv 1\bmod 4$.
@@
[…]
@@
For quadratic forms of rank six, Proposition 5.4 shows that invariants of the projective equivalence class are the discriminant $d(q)$ and $\epsilon_{p}(q)$ for all primes $p$ such that d(q)-d(q) is a square in $\mathbb{Q}_{p}$.
@@
    $H$ is $C_{4}$ or $D_{8}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\bmod 4$ and d(q)-d(q) is a square in $\mathbb{Q}_{p}$,
@@
    $H$ is $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$ and $\epsilon_{p}(q)=1$ for all primes $p$ such that $p\equiv 1\mod 3$ and d(q)-d(q) is a square in $\mathbb{Q}_{p}$, or
@@
    $H$ is $A_{4}$, and $\epsilon_{p}(q)=1$ for all primes $p$ such that d(q)-d(q) is a square in $\mathbb{Q}_{p}$.
@@
  • If $H$ is isomorphic to $C_{4}$ or $D_{8}$, the the representation $\rho_{1}$ must contain an element of order 4. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent to a quadratic form $\langle a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies d(q)=bcd(q)=-bc and one can calculate its Hasse-Witt invariants to be
@@
    Thus, we see that $\epsilon_{p}(q)=1$ whenever both d(q)-d(q) and 1-1 are squares in $\mathbb{Q}_{p}$. However 1-1 is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 4$.
    Conversely, any form $q$ of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ for all $p$ such that d(q)-d(q) and 1-1 are both squares in $\mathbb{Q}_{p}$ is equivalent to one of the form $\langle a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Proposition 5.1 allows us to pick $\alpha\in\mathbb{Q}_{>0}$ such that $(\alpha,-1)_{p}=\epsilon_{p}(q)$ for all primes where d(q)-d(q) is a square in $\mathbb{Q}_{p}$. Thus we see that $\epsilon_{p}(q)(\alpha,-1)_{p}=1$ for all primes where d(q)-d(q) is a square in $\mathbb{Q}_{p}$. This allows us to apply Proposition 5.1 again to obtain $\beta\in\mathbb{Q}_{>0}$ such that $(\beta,-d(q))_{p}=\epsilon_{p}(q)(\alpha,-1)_{p}$ for all primes. Consequently we have $(\beta d(q),-d(q))_{p}=\epsilon_{p}(q)(\alpha,-1)_{p}$ for all primes. If we set $b=-\beta d(q)>0$, $a=\alpha b$ and $c=\beta$, then this yields a form equivalent to $q$.
  • If $H$ is isomorphic to $C_{3}$, $C_{6}$, $D_{6}$ or $D_{12}$, then $\rho_{1}$ must contain an element of order 3. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent to $\langle 3a,a,b,c,1,-1\rangle$, for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies d(q)=3bcd(q)=-3bc and
@@
    Thus, we see that $\epsilon_{p}(q)=(3,-1)_{p}$ whenever both d(q)-d(q) and 3-3 are squares in $\mathbb{Q}_{p}$. However 3-3 is a square in $\mathbb{Q}_{p}$ if and only if $p\equiv 1\bmod 3$. Since $(3,3)_{p}=-1$ if and only if $p=2,3$, this reduces to the condition that $\epsilon_{p}(q)=1$.
    As in the previous case a double application of Proposition 5.1 can be used to show that any form $q$ of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ for all $p$ such that d(q)-d(q) and 3-3 are both squares in $\mathbb{Q}_{p}$ is equivalent to $\langle 3a,a,b,c,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$.
Thus it remains to treat the case that $H$ is the alternating group $A_{4}$. Since we know the holonomy representation is faithful, we see that $\rho$ must contain a copy of the unique irreducible 3-dimensional representation of $A_{4}$. Thus the holonomy representation decomposes as $\rho\cong\rho_{1}\oplus\rho_{2}$, where $\rho_{1}$ is 3-dimensional and $\rho_{2}$ is 1-dimensional. By direct calculation one can show that every form in $\mathcal{Q}(\rho_{1})$ is equivalent to $\langle a,a,a\rangle$ for some $a\in\mathbb{Q}$. This means that $B$ appears in the commensurability class of $q$ if and only if $q$ is equivalent to $\langle a,a,a,b,1,-1\rangle$ for some $a,b,c\in\mathbb{Q}_{>0}$. Such a $q$ satisfies d(q)=abd(q)=-ab and
@@
Thus we have that $\epsilon_{p}(q)=1$ whenever d(q)-d(q) is a square in $\mathbb{Q}_{p}$. Once again, Proposition 5.1 can be used to obtain the converse. If $q$ is a quadratic form of signature $(5,1)$ such that $\epsilon_{p}(q)=1$ whenever d(q)-d(q) is a square in $\mathbb{Q}_{p}$, then there is $b\in\mathbb{Q}_{>0}$ such that $\epsilon_{p}(q)=(b,-d(q))_{p}=(bd(q),-d(q))_{p}$ for all $p$. If we set a=bd(q)a=-bd(q), then $q$ is equivalent to $\langle a,a,a,b,1,-1\rangle$. ∎